OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin2 x - 4 sin x + 1 = 0

OpenStudy (anonymous):

@keepitsimple @keeponbleeding

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@bridgetx516x

OpenStudy (anonymous):

4sin^2x - 4sinx+1=0 (2sin(x) -1)^2 = 0 sin(x) = 1/2 sin(x) = 1/2 x = pi/6 or 5pi/6

OpenStudy (anonymous):

Thank u soooo much!!!! @bridgetx516x

OpenStudy (anonymous):

sorry for not going through it step by step but i just am a little busy.. I wouldnt normally just give you the answer

OpenStudy (anonymous):

i understan

OpenStudy (anonymous):

@bridgetx516x